# Communication Cost Reduction for Subgraph Counting under Local Differential Privacy via Hash Functions

### v1

@@ VI-D1 Error Analysis for Various Download Costs
Figure 1 illustrates that our algorithm outperforms the leading-edge algorithm across all download costs. When the download cost is minimal, the enhancement in relative error can be as substantial as a 1000-fold reduction.
The performance gap between our algorithm and the one presented in [22] widens as the communication cost decreases. This is attributed to the sampling rate $\mu^{*}$ dependency. To recap, $\mu^{*}$ represents the equivalent download cost reduction, equating to $\mu_{c}/s^{2}$ for GRR with central server sampling, and $\mu^{2}$ for ARROne. In the case of [22], the $S_{2}$ term scales with $1/\mu^{*}$ and the $C_{4}$ term with $1/\sqrt{\mu^{*}}$. For our algorithm, when setting $s=(1/\mu^{*})^{1/3}$ and $\mu_{c}=(\mu^{*})^{1/3}$, the $S_{2}$ term scales with $(\mu^{*})^{-2/3}$ and the $C_{4}$ term with $(\mu^{*})^{-1/3}$. This scaling results in our algorithm being more effective at lower communication costs compared to the approach in [22].
@@
When the download cost reaches its maximum — a scenario that occurs when no sampling is implemented — our mechanism and the leading mechanism exhibit similar performance. This outcome is anticipated, as in this scenario, both approaches essentially align with each other and with the classical local differential privacy-based triangle counting algorithm, as described in [20].
@@
VI-D2 Error Analysis for Upload Costs
The second experiment involves a comparison between the following three methods:
  1. 1.
    The leading edge method detailed in [22]
  2. 2.
    Our GRR mechanism enhanced the clipping method
  3. 3.
    Our GRR mechanism enhanced the smooth sensitivity
In this experiment, we have opted not to implement the four-cycle trick or the central server sampling, as neither method contributes to the reduction of upload costs.
@@ VI-D2 Error Analysis for Upload Costs
We can notice from the results that the error of our GRR mechanism tends to plateau at higher upload costs. This phenomenon can be attributed to the behavior of the variance term associated with the publishing of the triangle count, which becomes significant for moderately sized graphs. Specifically, with minimal sampling, the number of edges in the obfuscated graph diminishes at a rate comparable to the increase of $\omega$. Consequently, the local sensitivity remains relatively constant. However, as the sampling extent increases, the edge count in the obfuscated graph becomes too sparse for the law of large numbers to be effective.
As the sampling increases, the sum $\sum\limits_{k:a^{\prime}_{k}=1}a^{(i)}_{j,h_{j}(k)}$ becomes a collection of random variables where the likelihood of each being 1 is considerably less than the inverse of the number of terms in the sum. Consequently, most often this sum will equal 0, with occasional instances of 1 or 2. However, the maximum value of this sum across all $j$ will almost invariably be 1 or 2. This leads to the upper bound on the local sensitivity, denoted as $UB_{\text{LS}}(a^{\prime})=\max\limits_{j}\left(\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{j,h_{j}(k)}+\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{k,h_{k}(j)}\right)$, being equivalent to either $\omega_{j}$ or $2\omega_{j}$. Given that $\omega_{j}$ scales linearly with the sampling factor, it follows that for high levels of sampling, the noise added to the count also increases almost linearly with the sampling factor.
VI-D3 Error Analysis for Various Graph Sizes and Privacy Budgets
@@
Figure 2: Comparison on relative error between our algorithm and the leading edge algorithm for various graph sizes generated from the Wikipedia Article Network graph with $\varepsilon=1$ and a download cost reduction of 1000. In the legend, the abbreviation ’CSS’ denotes our Central Server Sampling method.
Figure 3: Comparison on relative error between our algorithm and the leading edge algorithm for various privacy budgets on the Wikipedia Article Network with a download cost reduction of 1000. In the legend, the abbreviation ’CSS’ denotes our Central Server Sampling method.
@@
Figures 2 and 3 also showcase that our algorithm outshines the current state-of-the-art across a wide array of parameters. In these experiments, we tested various graph sizes and privacy budgets, observing similar trends for both algorithms. Notably, even though the performance gap between the two algorithms appears to expand as the privacy budget increases, the difference in performance consistently remains around three orders of magnitude, irrespective of the graph size.
VII Conclusion
@@
Our work introduces a private graph publishing mechanism that provides unbiased estimations of all graph edges while maintaining low communication costs. The effectiveness of this method stems from integrating two key components: the application of universal hashing functions to achieve uniform edge partitions and the amplification of budget efficiency through sampling within each group. This synergy leads to a significant reduction in download costs, quantifiable as a factor of $\mathcal{O}(s^{3})$, where $s$ represents the size of the groups.
@@
We then demonstrated its utility in accurately estimating the number of triangles in a graph. Furthermore, we elaborated on the application of smooth sensitivity to these problems, ensuring that the resulting estimations remain unbiased. Our experiments have demonstrated that these methods offer substantial improvements in precision over existing state-of-the-art approaches for the counting.
@@
Acknowledgment
Quentin Hillebrand is supported by KAKENHI Grant 20H05965. Vorapong Suppakitpaisarn is partially supported by KAKENHI Grant 21H05845 and 23H04377. Tetsuo Shibuya is partially supported by KAKENHI Grant 20H05967, 21H05052, and 23H03345.

### v2

@@ 7.1.1 Error Analysis for Various Download Costs
Figure 3 illustrates that our algorithm outperforms the leading-edge algorithm across all download costs. When the download cost is minimal, the enhancement in relative error can be as substantial as a 1000-fold reduction. As the cost of communication goes down, the difference in performance between our method and the one described in Imola et al. (2022a) becomes more noticeable.
Parameters selection for GroupRR
Recall that the download cost of the randomized response algorithm is approximately $\mathsf{c}n^{2}$, for some constant $\mathsf{c}$ depending on the privacy budget $\varepsilon$. Suppose our budget for the download cost is $M$; then we aim to reduce the cost by a factor of $\mu^{*}=\mathsf{c}n^{2}/M$. As discussed in the previous section, for GroupRR with central server sampling, this reduction factor is given by $\mu^{*}=\mu_{c}/s^{2}$, where $\mu_{c}$ and $s$ are tunable parameters in GroupRR. To achieve the desired reduction, we set $s=(1/\mu^{*})^{1/3}$ and $\mu_{c}=(\mu^{*})^{1/3}$. Under this setting, the $S_{2}$ term in the error bound (as described in Theorem 11) scales as $(\mu^{*})^{-2/3}$, while the $C_{4}$ term scales as $(\mu^{*})^{-1/3}$.
Parameters selection for ARROne
On the other hand, in the context of ARROne, where $\mu$ represents the sampling rate of the mechanism, the reduction in download costs amounts to $\mu^{2}$. We hence set $\mu$ to $\sqrt{\mu^{*}}$. By that, the $S_{2}$ factor in the error term scales with $1/\mu^{*}$, and the $C_{4}$ factor scales with $1/\sqrt{\mu^{*}}$. This scaling behavior highlights that our algorithm proves more effective at lower communication costs compared to the method proposed in Imola et al. (2022a).
@@
When the download cost reaches its maximum — a scenario that occurs when no sampling is implemented — our mechanism and the leading mechanism exhibit similar performance. This outcome is anticipated, as in this scenario, both approaches essentially align with each other and with the classical local differential privacy-based triangle counting algorithm, as described in Imola et al. (2021).
@@
Figure 4: Comparative analysis of relative error between our algorithm and ARR across different download costs applied to the Google+ graph, assuming a privacy budget of $\varepsilon=1$. The values of the parameter $s$ used for this experiment vary from 1 to 32, resulting in values of $\mu_{c}$ varying from 1 to $32,768$. The term ‘CSS’ in the legend stands for our central server sampling method.
Figure 4 illustrates that the performance of our method compares to the state of the art becomes better as the size of the graph increases. Indeed, as the download cost needs to be reduced by a larger factor to keep the same communication, the gap in relative errors becomes bigger. For this experience we only used the clipping version of our algorithm as smooth sensitivity is slow for large graphs.
7.1.2 Error Analysis for Upload Costs
Throughout this section, we fix the download-reduction target at $\mu^{*}=1/1000$. As discussed in Section 7.1.1, this choice corresponds to GroupRR parameters $s=10$ and $\mu_{C}=1/10$. For ARROne, we set $\mu\approx 1/31.6$ (i.e., $\mu\approx 0.0316$).
@@ 7.1.2 Error Analysis for Upload Costs
7.1.3 Error Analysis for Various Graph Sizes and Privacy Budgets
(a) Wikipedia Article Network
(b) Facebook
@@ 7.1.3 Error Analysis for Various Graph Sizes and Privacy Budgets
Figure 5: Comparing the relative error between our algorithm and the state-of-the-art approach (left plots) across different graph sizes $(n)$, derived from the Wikipedia Article Network and Facebook graphs with $\varepsilon=1$ (right plots) under varying privacy budgets. In both experiments, we maintain a download reduction setting of $\mu^{*}=1000$. Note that in the legend, ‘CSS’ represents our central server sampling method.
@@
Figure 5 shows that our algorithm improves the current state-of-the-art across a wide array of parameters. In these experiments, we tested various graph sizes and privacy budgets, observing similar trends for both algorithms. Notably, even though the performance gap between the two algorithms appears to expand as the privacy budget increases, the difference in performance consistently remains around three orders of magnitude, irrespective of the graph size.
7.1.4 Runtime for Various Graph Sizes
(a) Wikipedia Article Network
(b) Facebook
Figure 6: Comparison on runtime between our algorithm and the leading edge algorithm for various graph sizes generated from the Wikipedia Article Network and Facebook graphs with $\varepsilon=1$ and a download cost reduction of 1000. In the legend, the abbreviation ‘CSS’ denotes our central server sampling method.
Figure 6 presents the execution times for the state-of-the-art ARR algorithm, compared to two variants of our GroupRR algorithm: one with clipping and another with smooth sensitivity, across different graph sizes. The results indicate that GroupRR with smooth sensitivity incurs significantly longer execution times, whereas GroupRR with clipping shows running times comparable to ARR. This discrepancy in performance can be attributed to their asymptotic complexities: both ARR and GroupRR with clipping operate in $\mathcal{O}\left(d_{max}m\right)$, whereas the complexity for GroupRR with smooth sensitivity is $\mathcal{O}\left(nm\right)$. The graphs, plotted on a logarithmic scale, do not form perfect straight lines, reflecting that in our model for generating graphs of varying sizes, the average degree does not scale proportionally with graph size $n$.
This experiment demonstrates the balance between speed and accuracy required when choosing between our two methods. GroupRR with clipping offers good accuracy with an execution time comparable to state-of-the-art methods. Conversely, GroupRR with smooth sensitivity prioritizes accuracy but at the expense of execution speed. This trade-off becomes particularly significant when $d_{max}$, the maximum degree, is small relative to the number of users $n$.
7.1.5 Error Analysis on Power-law Graphs
Figure 7: Comparative analysis of relative error between our algorithm and ARR across different graph sizes on synthetically generated power-law graphs, assuming a privacy budget of $\varepsilon=1$. For this experiment, we set the parameter $s$ to $5$. The term ‘CSS’ in the legend stands for our central server sampling method.
Figure 7 shows the relative error of our algorithms as well as the state-of-the-art ARROne for various graph sizes. Each graph is randomly samples such that the degree distribution follows a power law of exponent 2. This means that the expected number of nodes with a degree $d$ is proportional to $d^{-2}$.
Consequently, the graph’s maximum degree is approximately one-third of the number of nodes. Even in this regime—where the bounded-degree condition $d_{i}<m\approx n/s$ required for our theoretical guarantees is violated—our algorithms still outperform the state-of-the-art ARROne.
7.2 4-cycles Counting
To validate the versatility of our framework, we extended our experiments to include counting cycles of length 4, or 4-cycles, utilizing the same framework employed in our triangle counting methodology. This method is described in Eden et al. (2023), Hillebrand et al. (2025) and Suppakitpaisarn et al. (2025). The number of 4-cycles that include edges $(i,i^{\prime})$ and $(i,i^{\prime\prime})$ can be estimated by $\sum_{j}\tilde{a}(i^{\prime},j)\tilde{a}(j,i^{\prime\prime})$ and the number of 4-cycle that the node $i$ participates in is given by $\sum_{i^{\prime},i^{\prime\prime}:(i,i^{\prime}),(i,i^{\prime\prime})\in E}\sum_{j}\tilde{a}(i^{\prime},j)\tilde{a}(j,i^{\prime\prime})$.
Given the larger size of the subgraphs of interest, we opted to use smaller graphs for our experiments to effectively manage complexity and computational demands. To establish the robustness and applicability of our results, we conducted our 4-cycles counting experiments on two specific graphs. The first is the Twitter Interaction Network for the US Congress (Fink et al., 2023), which represents Twitter interactions of the 117th United States Congress, comprising 475 nodes and 13,289 edges. The second graph is the email-Eu-core network (Leskovec et al., 2007; Yin et al., 2017), derived from email interactions among members of a major European research institution, containing 1005 nodes and 25,571 edges. These selections allow us to provide substantive evidence of the generality of our findings across different types of networks.
We assessed accuracy using the $\ell_{2}$-error computed as $\sqrt{\sum\limits_{t=1}^{10}(\hat{x}_{t}-x)^{2}}$, where $x$ is the ground truth and $\hat{x}_{t}$ is the estimated value obtained in trial $1\leq t\leq 10$.
Figure 8: Comparison on $\ell_{2}$-error between our algorithm and ARR for various sampling factors on the Twitter Interaction Network for the US Congress and the email-Eu-core network with $\varepsilon=1$.
The experimental outcomes are displayed in Figure 8. It is evident that the error patterns are consistent across both datasets. With either method, regardless of the sampling factor used, the error is greater in the email-EU-core dataset. This aligns with expectations since the $\ell_{2}$-error is a property that tends to increase with the graph’s size.
Nonetheless, the rate at which the discrepancy between the two methods grows is the same across both datasets. When the full dataset is used—i.e., without any communication reduction—the performance of ARR and GroupRR is comparable and aligns with that of the standard algorithm without communication constraints (Imola et al., 2021). However, as shown in the graphs, under a sampling factor of $100$, the error associated with ARR increases by approximately a factor of $10^{3}$ relative to GroupRR. According to Theorem 4, the $\ell_{2}$-error in estimating a single edge scales linearly with the sampling factor $s$ for GroupRR, but quadratically (i.e., as $s^{2}$) for ARR. Since each $4$-cycle estimation involves two edge estimates, the overall $\ell_{2}$-error scales as $s^{2}$ for GroupRR and as $s^{4}$ for ARR. This implies a theoretical error gap of $s^{2}=10^{4}$ between the two methods. The observed reduction of this gap to $10^{3}$ is attributed to the fact that, in the absence of sampling, ARR achieves slightly better accuracy than GroupRR.
8 Conclusion
@@
Our work introduces a private graph publishing mechanism that provides unbiased estimations of all graph edges while maintaining low communication costs. The effectiveness of this method stems from integrating two key components: the application of linear congruence hashing to achieve uniform edge partitions and the amplification of budget efficiency through sampling within each group. This synergy leads to a significant reduction in download costs, quantifiable as a factor of $\mathcal{O}(s^{3})$, where $s$ represents the size of the groups.
@@
We then demonstrated its utility in accurately estimating the count of triangles and 4-cycles. Furthermore, we elaborated on the application of smooth sensitivity to these problems, ensuring that the resulting estimations remain unbiased. Our experiments have demonstrated that these methods offer substantial improvements in precision over existing state-of-the-art approaches for the countings.
@@
9 Acknowledgments
Quentin Hillebrand is partially supported by KAKENHI Grant 20H05965, and by JST SPRING Grant Number JPMJSP2108. Vorapong Suppakitpaisarn is partially supported by KAKENHI Grant 21H05845 and 23H04377. Tetsuo Shibuya is partially supported by KAKENHI Grant 20H05967, 21H05052, and 23H03345. The authors wish to express their thanks to the anonymous reviewers whose valuable feedback greatly enhanced the quality of this paper.